Hi I need the steps for this
P = 2W + 2L.
To solve for " W " in this formula, which steps should be followed?
Thanks :)
: Hi I need the steps for this
P = 2W + 2L.
To solve for " W " in this formula, which steps should be followed?
Thanks :)

Answer by nerdybill(1052)

To solve for W, we need to isolate it to one side of the equation -- doesn't matter which side.
.
P = 2W + 2L
Start by subtracting 2L from both sides:
P - 2L = 2W
Finally, divide both sides by 2:
(P-2L)/2 = W
or, you could write:
P/2 - L = W